Sally has 52 coins all dimes and nickels that have the total value of $4.20. How many nickels does this bitch have?
Answer by Alan3354(69443)

Sally has 52 coins all dimes and nickels that have the total value of $4.20. How many nickels does this bitch have?
=================
n + d = 52
5n +10d = 420
--------------------
Or:
52 nickels = 260
Each nickel replaced by a dime adds 5 cents.
420 - 260 = 160
160/5 = 32 replacements by dimes.